OpenStudy (anonymous):

Verify that f and g are inverse functions f(x)=5x^2-2, x greater than/equal to 0; g(x)= (x+2 over 5) ^1/2

OpenStudy (anonymous):

@ash2326 @ganeshie8

OpenStudy (superdavesuper):

u can prove that by assuming y=f(x)=5x^2-2 then rearrange so that u get the eqn in the form of x= a function of y. i will do the first step here: y=5x^2-2 so 5x^2=y+2...simplify it further so the left side is x plz

OpenStudy (anonymous):

sqrt (y+2 divided by 5) =x @superdavesuper

OpenStudy (superdavesuper):

yup yup then substitute the new eqn: replace y with x and x with y....u will get....

OpenStudy (anonymous):

so sqrt (x+2 divided by 5) =y

OpenStudy (superdavesuper):

and viola heres ur proof! :)

OpenStudy (anonymous):

what do you do with g(x)= (x+2 over 5) ^1/2?

OpenStudy (anonymous):

@superdavesuper

OpenStudy (superdavesuper):

ur Q asks u to verify f(x) n g(x) are inverse function - u just did it

OpenStudy (anonymous):

oh okay and they arent since f(x) doesnt have ^1/2

OpenStudy (superdavesuper):

sqrt( ) = ( )^1/2

OpenStudy (superdavesuper):

so they are the same :)

OpenStudy (anonymous):

gotcha thanks :)

OpenStudy (superdavesuper):

welcome n thanks for the medal!